Use differentials to approximate the given value. 5.8 2) 10 points
Answer #1

5.82 Letf(x) =x2 , x = 6·Δι = 5.8-6 =-02. Av = (x+Δι)2-r = 5.82-62 = 5.82-36 5.82 = 36+4. dv ах -2x6x(-0.2) From (1) 5.82 = 3